RAVEN69:

What is x 3+y 3+z 3=k

Nina001:

easy

RAVEN69:

I can;t figure it out I need help

Nina001:

I'm gonna help you

RAVEN69:

Thank you

Nina001:

\(x=\sqrt[3]{k-y^{3}-z^{3}}\)

RAVEN69:

My teacher isn't good at helping so thank you

Nina001:

lemme break it down for you

RAVEN69:

Yes please

Nina001:

Step \(x^{3}+y^{3}+z^{3}=k\)\(\xpmhighlightbox{bgcolor=#3A3F50, underlinecolor=#3A3F50}{x^{3}+y^{3}+z^{3}}=k\)Move the variables to the right-hand side and change their signs \(x^{3}=\xpmhighlightbox{bgcolor=#3A3F50, underlinecolor=#3A3F50}{-y^{3}-z^{3}}+k\)Step \(x^{3}=-y^{3}-z^{3}+k\)\(\xpmhighlightbox{bgcolor=#3A3F50, underlinecolor=#3A3F50}{x^{3}}=\xpmhighlightbox{bgcolor=#3A3F50, underlinecolor=#3A3F50}{-y^{3}-z^{3}+k}\)Take the root of both sides of the equation \(\xpmhighlightbox{bgcolor=#3A3F50, underlinecolor=#3A3F50}{x}=\xpmhighlightbox{bgcolor=#3A3F50, underlinecolor=#3A3F50}{\sqrt[3]{-y^{3}-z^{3}+k}}\)Step \(x=\sqrt[3]{-y^{3}-z^{3}+k}\)\(x=\sqrt[3]{\xpmhighlightbox{bgcolor=#3A3F50, underlinecolor=#3A3F50}{-y^{3}-z^{3}+k}}\)Use the commutative property to reorder the terms\(x=\sqrt[3]{\xpmhighlightbox{bgcolor=#3A3F50, underlinecolor=#3A3F50}{k-y^{3}-z^{3}}}\)Solution\(x=\sqrt[3]{k-y^{3}-z^{3}}\)

Nina001:

see

Nina001:

I'm so smarttt

RAVEN69:

Thank you Nina

Nina001:

you are welcome

Nina001:

bro I just heard a firework..

Nina001:

they buggin

RAVEN69:

Oh I'm scared for you

Nina001:

@raven69 wrote:
Oh I'm scared for you
nah word I'm scared for my self

RAVEN69:

IT'S NOT EVEN DARK YET

Nina001:

FOR REAL

Nina001:

they not even gonna be able 2 see shi